Hello, I could reeally use some help with this one. Thanks! The price of one brand of athletic shoes is $5 more than another brand. If the prices of the two pairs of shoes add to $230, how much does each pair cost?
Answer by tutorcecilia(2152)

Shoe A=(Shoe B + $5.00)
Shoe B = (B)
.
Shoe A + Shoe B =$230
(B+5)+(B)=230 [Simplify]
B+5+B=230 [Solve for B]
2B+5=230
2B=230-5
2B=225
B=$112.50 (Cost of Shoe B)
.
Therefore Shoe A = (Shoe B + $5)=112.50+5=$117.50 (Cost of Shoe A)
$112.50+$117.50=$230.00